# Count-mean Sketch as an Optimized Framework for Frequency Estimation with Local Differential Privacy

### v1

@@ 3.5 Unbiased CMS with Imperfect Hashing
The condition of these two corollaries is independent of $\epsilon$. If $\tau=0.01$, we have $q>4.5$ to satisfy the two corollaries above.
On the other hand, one may prefer using the original frequency estimator $\hat{f}(x)$ in Eq. (2) instead of $g(x)$, while still employing the imperfect hashing. The following theorem will demonstrate how large $q$ is required to make the deviation from $MSE(\hat{f}(x))$ negligible.
Theorem 3.9.
When
$\frac{1}{4(m-1)q(q+1)}\leq\sqrt{\tau Var(\hat{f}(x))}$
and
$\frac{8q^{2}+8q+1}{16q^{2}(q+1)^{2}}\leq\tau,$
we have
$MSE(f(x)|\mathcal{H}^{\prime})<(1+2\tau)MSE(\hat{f}(x)),$
where $MSE(\hat{f}(x))$ represents the MSE of the $f(x)$ with the perfect hashing.
Notably, $Var(\hat{f}(x))=\Theta(1/n)$, so $q=\Theta(n^{1/4})$. That is, the size of $\mathcal{H}^{\prime}$ has to increase with the size of the dataset.
All the theorems and corollaries in this section are proven in Appendix E.
3.6 Correcting the original CMS
@@
In the original CMS paper [7], CMS randomly generates $k$ hash functions, and each hash function $h_{j}(x)$ is generated by sampling $d$ mutually independent random variables, each uniformly distributed among $[m]$. Subsequently, $\forall(x,x^{\prime})\in[d]^{2},c_{j}(x,x^{\prime})$ can be considered as a Bernoulli random variable with a probability of $\frac{1}{m}$ of being one. Since the $k$ hash functions are generated independently, $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}$ is equivalent to a binomial random variable, whose expectation is $\frac{1}{m}$. However, the probability of the output of a binomial distribution exactly equaling its expectation is very low, so $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}$ will likely deviate from $\frac{1}{m}$ to some extent (see Slud’s Inequality [19]). Thus, $E[\hat{f}(x)]$ will likely deviate from $f(x)$ to some extent too, i.e., $\hat{f}(x)$ has bias. Notably, this bias is independent of the size of the dataset, so it will still exist even when an infinite amount of data is given for a single run of CMS. This conclusion is also applicable to the non-private Count Sketch families, which are also proven to have bias [20] if they query a dataset only once.
@@
The proof is presented in Appendix F.1. The variance indicates that the bias is $O(\frac{1}{\sqrt{(m-1)k}})$. However, the original CMS paper [7] derived $E[\hat{f}(x)]$ to be unbiased. We found that they treated $c_{j}(x,x^{\prime})$ as independent random variables for every pair of $X^{(i)}$ and $X^{(i^{\prime})}$, which is incorrect because all pairs with the same values $x$ and $x^{\prime}$ will always have the same $c_{j}(x,x^{\prime})$.
Eq. (7) demonstrates how the bias of $\hat{f}(x)$ contributes to its MSE. The following theorem demonstrates how to prevent the bias from being a dominant term in $MSE(\hat{f}(x))$:
Theorem 3.11.
If
$\frac{\delta\tau Var(R|\neq)}{n}>\frac{1}{(m-1)k}$
is satisfied, where $\tau$ is a small number, like 0.1, we have
$Pr\biggl{(}(E[\hat{f}(x)]-f(x))^{2}\leq\tau Var(\hat{f}(x))\biggr{)}\geq 1-\delta.$ (22)
The proof is presented in Appendix F.2. This implies that the number of hash functions $k$ should be $\Omega(\frac{n}{m})$ to ensure that the bias of $f(x)$ is negligible. However, this constraint was not mentioned in the original paper.
The original paper also has an issue in its variance calculation, where they required the hashing family $\mathcal{H}$ to be 3-wise-independent because they assumed any $X^{(i_{1})}$ and $X^{(i_{2})}$ were correlated, which was problematic. In contrast, the non-private Count Sketch [20] only requires pairwise independence, which is consistent with our work. Notably, the LDP perturbation only affects the accuracy after the hashing, so it should not alter the pairwise-independence requirement for the hashing family.
4 Connection with Existing Frequency Estimation Algorithms
This section demonstrates that various frequency estimation algorithms proposed in the previous literature are equivalent or closely connected to CMS. First, we define the Transformation-and-Sampling Framework (TnS Framework) as follows:
@@ 4 Connection with Existing Frequency Estimation Algorithms
1. Transform the original value $x$ to a vector of elements $M[x]$, i.e., a row of a matrix $M$.
2. Uniformly sample an element from $M[x]$, denoted as $M[x][j]$
3. Perturb the sampled element $M[x][j]$ with some LDP mechanism.
The above process is equivalent to (1) uniformly sampling a column from $M$, denoted as $M[:,j]$, (2) mapping the original value $x$ to $M[x,j]$, and (3) perturbing it. The column here functions as a hash function, and all the columns of $M$ comprise a hashing family. Thus, we have proved:
@@ 4.1 Hadamard Encoding
It is evident that HE belongs to the TnS framework, so it is equivalent to CMS. Additionally, Appendix G proves that decoding process of HE is equivalent to Eq. (2). Moreover, all the rows of a Hadamard matrix, excluding the first element, comprise a 3-wise-independent hashing family [21], mapping to either -1 or 1. Therefore, HE is equivalent to the CMS+RR using pairwise-independent hashing with $m=2$.
@@ 4.3 Recursive Hadamard Response
Recursive Hadamard Response (RHR) [12] is an improvement of Hadamard Encoding, increasing accuracy by communicating more bits. It (1) splits the original value $x$ into $x//2^{b-1}$ (floor division) and $x\bmod 2^{b-1}$; (2) creates a Hadamard matrix with a size larger than $d//2^{b-1}$, (3) select a row from the Hadamard matrix with an index equal to $(x//2^{b-1})+1$, and (4) uniformly samples an element from the selected row as $s$, and (5) perturbs $s\times(x\bmod 2^{b-1})$ using randomized response.
Steps 2 - 5 can be fitted in the TnS framework, mapping $x$ to $2^{b}$ possible values. However, the collision probability of two inputs is not uniform. If two inputs $x_{1}$ and $x_{2}$ have the same mod (i.e., $x_{1}\bmod 2^{b-1}=x_{2}\bmod 2^{b-1}$), their collision probability is $1/2$. Otherwise, the probability is zero. In CMS, any pair of inputs have the same collision probability of $1/m$, which is $2^{-b}$ given $2^{b}$ hashed values. Thus, RHR is a skewed version of CMS.
If the dataset is randomly generated, i.e., every pair of objects may have the same mod with a probability of $2^{b-1}$, then the hash collision probability becomes $2^{-b}$. Note that [12] sets $b=\epsilon$ if the dictionary is large enough. This is equivalent to $m=e^{\epsilon}$ in CMS, which is close but not identical to the optimized $m$ used by OCMS, so the accuracy of RHR at most has the same order as OCMS.
@@ 4.4 Local Hashing
Local Hashing [9] asks each client to (1) publicly and randomly generate a hashing function mapping $[d]$ to $[m]$, (2) hash the original value, and (3) perturb the hashed value using randomized response. It is straightforward to see that randomly generating a hashing function is equivalent to randomly selecting a hash function from the hashing family in CMS. Also, its decoding process is identical to Eq. (2). Therefore, this algorithm is equivalent to CMS. However, [9] did not compute the optimized $\widehat{MSE}$ correctly as they only considered $f(x)=0$, but Section 3.4 demonstrates that both $f(x)=0$ and $f(x)=1$ could be the maximum points of the MSE. Moreover, they did not optimize the communication cost of sending a hash function to the server, which will be discussed in Section 5.1.
@@ 4.5 Recursive CMS
Here, we assume all the CMS involved are unbiased, i.e., the hashing family is pairwise independent. Suppose the first CMS hashes the original values to $[m_{1}]$, and its perturbation algorithm is another CMS hashing its input to $[m_{2}]$. Given the pairwise independence, the collision probability $\bar{c}(x,x^{\prime})$ of two distinct input values is $\frac{m_{1}+m_{2}-1}{m_{1}m_{2}}$. Therefore, the recursive CMS is equivalent to a regular CMS with $m=\frac{m_{1}m_{2}}{m_{1}+m_{2}-1}$.
@@ 4.6 RAPPOR
The encoding of RAPPOR is equivalent to CMS if each object in RAPPOR belongs to its own cohort and there is only one hash function in each cohort. Assuming there is sufficient data and we can ignore the uncertainty from the LDP mechanism and hash-function assignment, the encoding of RAPPOR and CMS can be considered a linear transformation from $\mathcal{R}^{d}$ (frequencies of original values) to $\mathcal{R}^{km}$ (frequencies of the hashed values). For the unbiased CMS (Corollary 3.1), we have $km=d^{2}m>d$, so CMS can derive the original frequencies without bias. However, RAPPOR utilizes gradient descent to find the original frequencies and does not require $km\geq d$. [15] even chooses small $km$ on purpose to reduce computation. Unfortunately, this is problematic because multiple values in a high-dimension simplex will map to the same value in a low-dimension simplex, and the value RAPPOR finds may not be the original value.
@@ 5 Discussion
5.1 Accuracy, Communication Cost, and Optimality
Tables I and II compare the accuracy and communication cost among various frequency estimation algorithms, respectively. We will demonstrate that the optimized CMS is also the optimal algorithm when considering both accuracy and communication cost.
@@ 5.1 Accuracy, Communication Cost, and Optimality
OCMS $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
@@
SS $\sqrt{\frac{8}{\pi}}\frac{de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
@@
s-RP $\frac{de^{\epsilon/4}}{e^{\epsilon/2}-1}$ $\frac{de^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
TABLE I: Accuracy comparison among frequency-estimation algorithms
  • Note: HE, RHR, OLH, OCMS, CMS+HE, SS, a-RP, and s-RP represent Hadamard Encoding, Recursive Hadamard Response, Optimized Local Hashing, Optimized CMS, CMS with Hadamard Encoding, Subset Selection, asymmetric RAPPOR, and symmetric RAPPOR, respectively. Here, OCMS represents MSE-OCMS and $l$-OCMS when comparing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively. The data source is presented in Section H.
@@
MSE-OCMS $2\log d+\frac{\epsilon}{2}$
$l$-OCMS $2\log d+\epsilon$
@@
  • Note: see Table I for the meaning of abbreviations. The communication cost only accounts for the cost between the server and one client. The data source is presented in Section H.
Table I shows that the optimized CMS (OCMS) is optimal in $l_{2}$ loss and the worst-case MSE, even considering the constant term. Its $l_{1}$ loss is order-optimal. Since HE, OLH, and CMS+HE are equivalent to unbiased CMSs without the optimized parameters, it is expected that they are not optimal in at least one dimension. Only OCMS and s-RP are optimal in the worst-case MSE, but Table II shows that s-RP requires $d$ bits to communicate while OCMS only requires $2\log_{2}d+\frac{\epsilon}{2}$ bits. If the communication cost is required to be $O(\log d)$, only RHR has the same order of accuracy as OCMS in $l_{1}$ and $l_{2}$ losses. However, empirical study (Section 6) demonstrates that OCMS outperforms RHR in all datasets, consistent with the theoretical analysis in Section 4.3. Therefore, OCMS is the only algorithm optimized for the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss while communicating efficiently.

### v2

@@ 3.5 Unbiased CMS with Imperfect Hashing
for MSE-OCMS+RR and $l$-OCMS+RR, respectively.
The theorem and corollary are proven in Appendix H. Notably, the condition of this corollary is independent of $\epsilon$. If $\tau=0.01$, we have $q>4.5$ to satisfy the two corollaries above. Therefore, one could use $g(x)$ to approximate $\hat{f}(x)$ if $p^{l}\geq 5m$. Since $a_{0}$ and $a_{1}$ of $\mathcal{H}_{api}$ can be any numbers in $\mathbb{F}(p^{l})$, we have $|\mathcal{H}_{api}|=(p^{l})^{2}=p^{2\lceil\log_{p}\max\{d+1,5m\}\rceil}$, and $2\lceil\log_{2}\max\{d+1,5m\}\rceil$ bits are sufficient to represent a hash function in $\mathcal{H}_{api}$ if $p=2$.
3.6 Implementation
Algorithm 1 OCMS+RR (Part 1)
1:$d$: dictionary size
2:$\epsilon$: privacy factor
3:Mode := either MSE-OCMS+RR or $l$-OCMS+RR
4:$f^{*}$: prior knowledge on the upper bound of $f(x)$. Default value is one.
[…]
6:
7:function FiniteFieldSize($d$, $m$)
8:     $v\coloneq\max\{d+1,5m\}$ // see Corollary 3.5 for why $5m$
[…]
11:     end if
12:     return $2^{\lceil\log_{2}(v)\rceil}$
13:end function
14:
15:function HashRange($\epsilon,d,f^{*}$, Mode)
16:     if Mode is MSE-OCMS+RR then
17:         if $f^{*}\geq\frac{1}{2}$ then
18:              return round( $e^{\epsilon/2}+1$ )
19:         end if
20:         $\Delta\coloneq e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]}$
21:         return round( $\frac{\Delta}{f^{*}e^{\epsilon}+1-f^{*}}+1$ )
22:     else if Mode is $l$-OCMS+RR then
23:         $\Delta\coloneq e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}$
24:         return round( $\frac{\Delta}{e^{\epsilon}+d-1}+1$ )
25:     end if
26:end function
27:
28:$m\coloneq$ :- HashRange($\epsilon,d,f^{*}$, Mode)
29:$p_{\mathcal{H}}\coloneq$ :- FiniteFieldSize($d,m$) // i.e., $p^{l}$ in Section 3.5.
30:
31:function Hash($a_{0},a_{1},x,m,p_{\mathcal{H}}$)
[…]
33:         $v\coloneq(a_{0}+a_{1}x)\mod{p_{\mathcal{H}}}$
34:     else
[…]
37:     end if
38:     return $v\mod{m}$
39:end function
40:
41:// return $Q[v,u]$ in Section 2.1.
42:function Decode($u$: perturbed value, $v$: value of interest, $m$, $\epsilon$)
43:     if $u=v$ then
44:         return $\frac{e^{\epsilon}+m-1}{e^{\epsilon}-1}$
45:     end if
46:     return $\frac{-1}{e^{\epsilon}-1}$
47:end function
48:
49:function Client($x$: true value of a client, $X^{(i)}$)
50:     Uniformly Sample $a_{0}$ and $a_{1}$ from $[0,p_{\mathcal{H}}-1]$.
51:     $y\coloneq$ :- Hash($a_{0},a_{1},x,m,p_{\mathcal{H}}$) // i.e., $h_{j^{(i)}}(X^{(i)})$
52:     return RandomizedResponse($y$, $m$, $\epsilon$), $a_{0},a_{1}$
53:end function
Algorithm 2 OCMS+RR (Part 2)
54:function Server($x$: value of interest, $\mathbf{C}$: all responses from the clients)
[…]
56:     for $z,a_{0},a_{1}$ in $\mathbf{C}$ do
57:         $h_{x}\coloneq$ :- Hash($a_{0},a_{1},x,m,p_{\mathcal{H}}$) // i.e., $h_{j^{(i)}}(x)$
58:         $\hat{y}\coloneq$ :- Decode($z$, $h_{x}$, $m$, $\epsilon$) // i.e., $\hat{y}^{(i)}[h_{j^{(i)}}(x)]$
59:         $\hat{y}_{tot}$ += $\hat{y}$
60:     end for
61:     $n\coloneq$ :- Length($\mathbf{C}$)
62:     $q,\;r\coloneq\lfloor\>p_{\mathcal{H}}/m\>\rfloor,\;\;p_{\mathcal{H}}\mod{m}$
63:     $m^{\prime}\coloneq\frac{(mq+r)^{2}}{(2q+1)r+mq^{2}}$
64:     return $\frac{m^{\prime}}{n(m^{\prime}-1)}\hat{y}_{tot}-\frac{1}{m^{\prime}-1}$
65:end function
OCMS+RR (Algorithm 1) is implemented to not only adapt the optimized hash range $m$ proposed in Section 3.4, but also the approximately pairwise independent hashing family $\mathcal{H}_{api}$ and the corresponding estimator $g(x)$ in Section 3.5. The Client function perturbs each object $X^{(i)}$, and their perturbed values, along with their hashing functions, are sent to the server. The server then calls Server to estimate the frequency of any $x$, i.e., $f(x)$. In particular, when the dictionary size $d$ is small, i.e., $d<2^{64}-59$, the algorithm constructs the hashing family with $p^{l}$ as a prime number. Otherwise, it computes the minimum $2^{l}>d$ as $p^{l}$. The source code of OCMS+RR is available at https://github.com/mingen-pan/CountMeanSketch).
3.7 Issues in the original CMS
Most frequency estimators ensure that the estimated frequency converges to the actual frequency as the dataset size increases, provided each object is queried once. However, this section demonstrates that the estimation from the original CMS exhibits a constant bias, even with an infinitely large dataset.
@@ 3.7 Issues in the original CMS
The original CMS randomly generates $k$ hash functions, and each hash function $h_{j}(x)$ is generated by sampling $d$ mutually independent random variables, each uniformly distributed among $[m]$. Subsequently, $\forall x,x^{\prime}\in[d]$ with $x\neq x^{\prime}$, $c_{j}(x,x^{\prime})$ can be considered as a Bernoulli random variable with a probability of $\frac{1}{m}$ of being one. Since the $k$ hash functions are generated independently, $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}$ is equivalent to a binomial random variable, whose expectation is $\frac{1}{m}$. However, the probability of the output of a binomial distribution exactly equaling its expectation is very low, so $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}$ will likely deviate from $\frac{1}{m}$ to some extent (see Slud’s Inequality [20]). Given Corollary 3.1, $E[\hat{f}(x)]$ will likely deviate from $f(x)$ to some extent too, i.e., $\hat{f}(x)$ has bias.
@@
The proof is presented in Appendix I. The variance indicates that the bias is $O(\frac{1}{\sqrt{(m-1)k}})$. However, the original CMS paper [7] derived $E[\hat{f}(x)]$ to be unbiased. We found that they treated $c_{j}(x,x^{\prime})$ as independent random variables for every pair of $X^{(i)}$ and $X^{(i^{\prime})}$, which is incorrect because all pairs with the same values $x$ and $x^{\prime}$ will always have the same $c_{j}(x,x^{\prime})$.
The original paper also has an issue in its variance calculation, where they required the hashing family $\mathcal{H}$ to be 3-wise-independent because they assumed any $X^{(i_{1})}$ and $X^{(i_{2})}$ were correlated, which was problematic. In contrast, the non-private Count Sketch [21] only requires pairwise independence, which is consistent with our work. Notably, the LDP perturbation only affects the precision after the hashing, so it should not alter the pairwise-independence requirement for the hashing family.
4 Connection with Existing Frequency Estimators
This section demonstrates that various frequency estimators proposed in the previous literature are equivalent or closely connected to CMS. First, we define the Transformation-and-Sampling Framework (TnS Framework) as follows:
@@ 4 Connection with Existing Frequency Estimators
1. Transform the original value $x$ to a vector of elements $M[x]$, i.e., a row of a matrix $M$, where the $j$-th element belongs to the finite set $\mathcal{V}_{j}$.
2. Uniformly sample an element from the vector $M[x]$, denoted as $M[x][j]$.
3. Perturb the sampled element $M[x][j]$ within $\mathcal{V}_{j}$ using some LDP mechanism.
The above process is equivalent to (1) uniformly sampling a column from $M$, denoted as $M[:,j]$, (2) mapping the original value $x$ to $M[x,j]$, and (3) perturbing it. The column here serves as a hash function, and all the columns of $M$ comprise a hashing family. Thus, we have proved:
@@ 4.1 Hadamard Encoding
It is evident that HE belongs to the TnS framework, so it is equivalent to CMS. Additionally, Appendix J proves that the decoding process of HE is equivalent to Eq. (2). Moreover, all the rows of a Hadamard matrix, excluding the first element, comprise a 3-wise-independent hashing family [22], mapping to either -1 or 1. Therefore, HE is equivalent to the CMS+RR using pairwise-independent hashing with $m=2$.
@@ 4.3 Recursive Hadamard Response
Recursive Hadamard Response (RHR) [12] is an improvement of Hadamard Encoding, increasing precision by communicating more bits. It (1) splits the original value $x$ into $x//2^{b-1}$ (floor division) and $x\bmod 2^{b-1}$; (2) creates a Hadamard matrix with a size larger than $d//2^{b-1}$, (3) select a row from the Hadamard matrix with an index equal to $(x//2^{b-1})+1$, and (4) uniformly samples an element from the selected row as $s$, and (5) perturbs $s\times(x\bmod 2^{b-1})$ using randomized response.
Steps 2 - 5 can be fitted in the TnS framework, mapping $x$ to $2^{b}$ possible values. However, the collision probability of two inputs is not uniform. If two inputs $x_{1}$ and $x_{2}$ have the same mod (i.e., $x_{1}\bmod 2^{b-1}=x_{2}\bmod 2^{b-1}$), their collision probability is $1/2$. Otherwise, the probability is zero. In CMS, any pair of inputs have the same collision probability of $1/m$, which is $2^{-b}$ given $2^{b}$ hashed values. Thus, RHR is a skewed version of CMS+RR.
If the dataset is randomly generated, i.e., every pair of objects may have the same mod with a probability of $2^{b-1}$, then the hash collision probability becomes $2^{-b}$. Note that [12] sets $b=\epsilon$ if the dictionary is large enough. This is equivalent to $m=e^{\epsilon}$ in CMS, which is close but not identical to the optimized $m$ used by OCMS+RR, so the precision of RHR at most has the same order as OCMS+RR.
@@ 4.4 Local Hashing
Local Hashing [9] asks each client to (1) publicly and randomly generate a hashing function mapping $[d]$ to $[m]$, (2) hash the original value, and (3) perturb the hashed value using randomized response. It is straightforward to see that randomly generating a hashing function is equivalent to randomly selecting a hash function from the hashing family in CMS. Also, its decoding process is identical to Eq. (2). Therefore, this algorithm is equivalent to CMS. The original work [9] assumed $f(x)\approx 0$ when optimizing the MSE of their algorithm, but Section 3.4 demonstrates that both $f(x)=0$ and $f(x)=1$ could be the maximum points of the MSE. Moreover, they did not optimize the communication cost of sending a hash function to the server, which will be discussed in Section 5.1.
@@ 4.5 Recursive CMS
Here, we assume all the CMS involved are unbiased, i.e., the hashing family is pairwise independent. Suppose the first CMS hashes the original values to $[m_{1}]$, and its perturbation algorithm is another CMS hashing its input to $[m_{2}]$. Given pairwise independence, the collision probability $\bar{c}(x,x^{\prime})$ of two distinct input values is $\frac{m_{1}+m_{2}-1}{m_{1}m_{2}}$. Therefore, the recursive CMS is equivalent to a regular CMS with $m=\frac{m_{1}m_{2}}{m_{1}+m_{2}-1}$.
@@ 4.6 RAPPOR
The encoding of RAPPOR [15] is equivalent to CMS if each object in RAPPOR belongs to its own cohort and there is only one hash function in each cohort. The hashing of RAPPOR and CMS can be considered a linear transformation from $\mathcal{R}^{d}$ (frequencies of original values) to $\mathcal{R}^{km}$ (frequencies of the hashed values). For an unbiased CMS (Corollary 3.1), we have $km=d^{2}m>d$, so CMS can derive the original frequencies without bias. However, RAPPOR utilizes gradient descent to find the original frequencies and does not require $km\geq d$. The original RAPPOR[15] even chooses small $km$ on purpose to reduce computation. Unfortunately, this is problematic because multiple values in a high-dimension simplex will map to the same value in a low-dimension simplex, and the value RAPPOR finds may not be the original value.
@@ 5 Discussion
5.1 Precision and Communication Cost
Tables I and II compare the precision and communication cost among various frequency estimators, respectively. We will demonstrate that OCMS+RR is the leading algorithm when considering both precision and communication cost.
@@ 5.1 Precision and Communication Cost
OCMS+RR $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
@@
SS $\frac{2de^{\epsilon/2}}{e^{\epsilon}-1}$ $\frac{4de^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ $(\frac{e^{\epsilon}+1}{e^{\epsilon}-1})^{2}$
@@
RP $\frac{de^{\epsilon/4}}{e^{\epsilon/2}-1}$ $\frac{de^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$ $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$
TABLE I: Precision comparison among frequency-estimation algorithms when $d\gg e^{\epsilon}$.
  • Note: HE, RHR, OLH, OCMS+RR, CMS+HE, SS, a-RP, and RP represent Hadamard Encoding, Recursive Hadamard Response, Optimized Local Hashing, Optimized CMS+RR, CMS with Hadamard Encoding, Subset Selection, asymmetric RAPPOR, and symmetric RAPPOR, respectively. Here, OCMS+RR represents MSE-OCMS+RR and $l$-OCMS+RR when comparing $\widehat{MSE}$ and $l_{1}$ / $l_{2}$ losses, respectively. The data source is presented in Appendix C.
  • *: the original value from [10] is problematic, and Appendix C recalculates the value as shown.
@@
MSE-OCMS+RR $\max\{2\log d+\frac{\epsilon}{2},\frac{3}{2}\epsilon+6\}$
$l$-OCMS+RR $\max\{2\log d+\epsilon,3\epsilon+6\}$
@@
  • Note: see Table I for the meaning of abbreviations. The communication cost only accounts for the cost between the server and one client. The data source is presented in Appendix C.
Table I shows that the OCMS+RR is leading in reducing $l_{1}$ loss, $l_{2}$ loss, and the worst-case MSE, even considering the constant term. Since HE, OLH, and CMS+HE are equivalent to unbiased CMS+RR without the optimized parameters, it is expected that they are not leading in at least one dimension. Only OCMS+RR and RP are leading in the worst-case MSE, but Table II shows that RP requires $d$ bits to communicate while OCMS+RR only requires $\max\{2\log d+\frac{\epsilon}{2},\frac{3}{2}\epsilon+6\}$ bits. If the communication cost is required to be $O(\log d)$, only RHR has the same order of precision as OCMS+RR in $l_{1}$ and $l_{2}$ losses. However, empirical study (Section 6) demonstrates that OCMS+RR outperforms RHR in all datasets, consistent with the theoretical analysis in Section 4.3. Therefore, OCMS+RR is the only leading algorithm for the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss while communicating efficiently.